OpenStudy (anonymous):

Subtract. 6/12 - 4/12 A. 5/6 B. 10/12 C. 1/12 D. 1/6

OpenStudy (michele_laino):

here we have to compute this: \[\frac{6}{{12}} - \frac{4}{{12}} = \frac{{6 - 4}}{{12}} = ...?\]

OpenStudy (anonymous):

2/12

OpenStudy (michele_laino):

correct! Now we have to simplify that fraction. In order to do that, we can divide numerator and denominator by \(2\): \[\frac{6}{{12}} - \frac{4}{{12}} = \frac{{6 - 4}}{{12}} = \frac{2}{{12}} = \frac{{2/2}}{{12/2}} = ...?\]

OpenStudy (anonymous):

1/6?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Thanks

OpenStudy (michele_laino):

:)